# Similarity-Navigated Conformal Prediction for Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks have achieved remarkable accuracy in semi-supervised node classification tasks. However, these results lack reliable uncertainty estimates. Conformal prediction methods provide a theoretical guarantee for node classification tasks, ensuring that the conformal prediction set contains the ground-truth label with a desired probability (e.g., 95%). In this paper, we empirically show that for each node, aggregating the non-conformity scores of nodes with the same label can improve the efficiency of conformal prediction sets. This observation motivates us to propose a novel algorithm named Similarity-Navigated Adaptive Prediction Sets (SNAPS), which aggregates the non-conformity scores based on feature similarity and structural neighborhood. The key idea behind SNAPS is that nodes with high feature similarity or direct connections tend to have the same label. By incorporating adaptive similar nodes information, SNAPS can generate compact prediction sets and increase the singleton hit ratio (correct prediction sets of size one). Moreover, we theoretically provide a finite-sample coverage guarantee of SNAPS. Extensive experiments demonstrate the superiority of SNAPS, improving the efficiency of prediction sets and singleton hit ratio while maintaining valid coverage.
@@ 1 Introduction
Graph Neural Networks (GNNs), which process graph-structured data by the message-passing manner (Kipf and Welling, 2017; Hamilton et al., 2017; Velickovic et al., 2018; Xu et al., 2019), have achieved remarkable accuracy in various high-stakes applications, e.g., drug discovery (Li et al., 2022), fraud detection (Liu et al., 2023) and traffic forecasting (Jiang and Luo, 2022), where any erroneous prediction can be costly and dangerous (Amodei et al., 2016; Gao et al., 2019). To improve the reliability of prediction results, many methods have been investigated to quantify the model uncertainty (Gal and Ghahramani, 2016; Guo et al., 2017b; Kendall and Gal, 2017; Wang et al., 2021; Hsu et al., 2022; Tang et al., 2024), while these methods lack theoretical guarantees of quantification. Conformal prediction (CP), on the other hand, offers a systematic approach to construct prediction sets that contain ground-truth labels with a desired coverage guarantee  (Vovk et al., 2005; Romano et al., 2020; Angelopoulos et al., 2021; Huang et al., 2023a).
@@
CP algorithms utilize non-conformity scores to measure dissimilarity between a new instance and the training instances. The lower the score of a new instance, the more likely it belongs to the same distribution space as the training instances, thereby included in the prediction set. To improve the efficiency of prediction sets for GNNs, DAPS (Zargarbashi et al., 2023) smooths node-wise non-conformity scores by incorporating neighborhood information based on the assumption of network homophily. Similar to DAPS, CF-GNN (Huang et al., 2023b) introduces a topology-aware output correction model that learns to update prediction and then produces more efficient prediction sets or intervals with the inefficiency as the optimization objective. However, they only consider structural neighbors and ignore the effect of other nodes that are far from the ego node. This motivates us to analyse the influence of global nodes on the size of prediction sets.
@@
In this work, we show that aggregating the information of global nodes with the same label as the ego node benefits the performance of CP methods. We provide an empirical analysis by randomly selecting nodes with the same label as the ego node from an oracle perspective, where the ground-truth labels of all nodes are known, and then aggregating their non-conformity scores into the ego node. The results indicate that aggregating scores of these nodes can significantly reduces the average size of prediction sets. This suggests that the information of nodes with the same label could correct the non-conformity scores, thereby prompting the efficiency of prediction sets. Detailed analysis is presented in Subsection 3.1. However, during the testing phase, the ground-truth label of every test node is unknown. Inspired by the analysis, our key idea is to accurately identify and select as many nodes with the same label as the ego node as possible and aggregate their non-conformity scores.
@@
To verify the effectiveness of our method, we conduct thorough empirical evaluations on 10 datasets, including both small datasets and large-scale datasets, e.g., OGBN Products (Bhatia et al., 2016). The results demonstrate that SNAPS not only achieves the pre-defined empirical marginal coverage but also achieves better performance over the compared methods. For example, on OGBN Products, our method reduces the average size of prediction sets from 14.92 of APS to 7.68. Moreover, we adapt SNAPS to image classification problems. The results demonstrate that SNAPS reduces the average size of prediction sets from 19.639 to 4.079 – only $\frac{1}{5}$ of the prediction set size from APS on ImageNet (Deng et al., 2009).
@@
  • We empirically explain that non-conformity scores of nodes with the same label as the ego node play a critical role in its non-conformity scores.
@@
  • Extensive experimental results demonstrate the effectiveness of our proposed method. We show that SNAPS not only maintains the pre-defined coverage, but also achieve great performance in efficiency and singleton hit ratio.
@@ 2 Preliminary
Adaptive Prediction Sets. In the APS method, the non-conformity scores are calculated by accumulating the softmax probabilities in descending order. Formally, given a data pair $(\boldsymbol{x},y)$ and a predicted probability estimator $\pi(\boldsymbol{x})_{y}$ for $(\boldsymbol{x},y)$, where $\pi(\boldsymbol{x})_{y}$ is the predicted probability for class $y$, the non-conformity scores can be computed by: $s(\boldsymbol{x},y)=\sum_{i=1}^{|\mathcal{Y}|}\pi(\boldsymbol{x})_{i}\mathbb{I}[\pi(\boldsymbol{x})_{i}>\pi(\boldsymbol{x})_{y}]+u\cdot\pi(\boldsymbol{x})_{y}$, where $u\in[0,1]$ is a uniformly distributed random variable. Then, the prediction set is constructed as $\mathcal{C}(\boldsymbol{x})=\{y|s(\boldsymbol{x},y)\leq\hat{q}\}$.
Evaluation Metrics. The goal is to improve the efficiency of conformal prediction sets as much as possible while maintaining the empirical marginal coverage guarantee. Given the testing nodes set $\mathcal{V}_{\text{test}}$, the efficiency is defined as the average size of prediction sets: $\mathrm{Size}:=\frac{1}{|\mathcal{V}_{\text{test}}|}\sum_{v_{i}\in\mathcal{V}_{\text{test}}}|\mathcal{C}(\boldsymbol{x}_{i})|.$ The smaller the size, the more efficient CP is. The empirical marginal coverage is defined as $\mathrm{Coverage}:=\frac{1}{|\mathcal{V}_{\text{test}}|}\sum_{v_{i}\in\mathcal{V}_{\text{test}}}\mathbb{I}[y_{i}\in\mathcal{C}(\boldsymbol{x}_{i})]$. Although efficiency is a common metric for evaluating CP, singleton hit ratio (SH), defined as the proportion of prediction sets of size one that contain the true label, is also important (Zargarbashi et al., 2023). The formula of SH is defined as: $\mathrm{SH}:=\frac{1}{|\mathcal{V}_{\text{test}}|}\sum_{v_{i}\in\mathcal{V}_{\text{test}}}\mathbb{I}[\mathcal{C}(\boldsymbol{x}_{i})=\{y_{i}\}]$.
@@ 3.2 Similarity-Navigated Adaptive Prediction Sets
$\mathrm{Sim}(i,j)=\frac{\boldsymbol{x}_{i}^{\top}\boldsymbol{x}_{j}}{\|\boldsymbol{x}_{i}\|_{2}\cdot\|\boldsymbol{x}_{j}\|_{2}}.$ (3)
@@
Figure 3: The overall framework of SNAPS. (1) Basic non-conformity score function. We first use basic non-conformity score functions, e.g., APS, to convert node embeddings into non-conformity scores. (2) SNAPS function. We then aggregate basic non-conformity scores of $k$-NN with feature similarity and one-hop structural neighbors to correct non-conformity scores of nodes. (3) Conformal Prediction. Finally, we use conformal prediction to generate prediction sets, significantly reducing their size compared to the basic score functions.
@@
where $s(\cdot,\cdot)$ is the basic non-conformity score function and $\hat{s}(\cdot,\cdot)$ is the SNAPS score function. Both $\lambda$ and $\mu$ are hyperparameters, which are used to measure the importance of three parts of non-conformity scores. The framework of SNAPS is shown in Figure 3 and pseudo-code is in Appendix B.
@@ 3.3 Theoretical Analysis
To deploy CP for graph-structured data, the only assumption we should satisfy is exchangeability, i.e., the joint distribution of calibration and testing sets remains unchanged under any permutation. Several studies have demonstrated that non-conformity scores based on the node embeddings obtained by any GNNs models are invariant to the permutation of nodes while permuting their edges correspondingly in the calibration and testing sets. This invariance arises because GNNs models and non-conformity score functions only use the structures and attributes in the graph, without dependence on the order of the nodes (Zargarbashi et al., 2023; Huang et al., 2023b). Under this condition, we prove that SNAPS non-conformity scores are still exchangeable.
@@
Assume that all of nodes aggregated by SNAPS are the same label as the ego node. Let $\boldsymbol{S}$ be APS scores of nodes, where $\boldsymbol{S}_{ui}\in[0,1]$ is the score corresponding to node $u$ with label $i$. Let $E_{k}[\boldsymbol{S}_{ui}]$ be the average of scores corresponding to label $i$ of nodes whose ground-truth labels are $k$ and $\eta$ be $1-\alpha$ quantile of basic non-conformity scores with a significance level $\alpha$. If $E_{k}[\boldsymbol{S}_{uk}]\ll\eta$ and $E_{k}[\boldsymbol{S}_{ui}]-\eta\geq-\Delta$, where $\Delta>0$ and is very small, then
$\mathbb{E}[|\mathcal{C}(\boldsymbol{x})|]\leq\mathbb{E}[|\mathcal{\tilde{C}}(\boldsymbol{x})|]$
where $\mathcal{C}(\cdot)$ and $\mathcal{\tilde{C}}(\cdot)$ represent the prediction set from APS score function and SNAPS score function, respectively.
In other words, SNAPS consistently generates a smaller prediction set than basic non-conformity scores functions and maintains the desired marginal coverage rate. It is obviously that we can’t ignore a very important thing, which is to select nodes with the same label as ego node as correctly as possible, otherwise it will lead to a decrease in the efficiency of SNAPS.
@@ 4.1 Experimental Settings
Datasets. We consider ten datasets in experiments, including the common citation graphs: CoraML (McCallum et al., 2000), PubMed (Namata et al., 2012), CiteSeer (Sen et al., 2008), CoraFull (Bojchevski and Günnemann, 2018), Coauthor Physics (Physics) and Coauthor CS (CS) (Shchur et al., 2019) and the co-purchase graphs: Amazon Photos (Photos) and Amazon Computers (Computers) (McAuley et al., 2015; Shchur et al., 2019). Moreover, we consider two large-scale graph datasets, i.e., OGBN Arxiv (Arxiv) (Wang et al., 2020) and OGBN Products (Products) (Bhatia et al., 2016). Particularly, for CoraFull which is highly class-imbalanced, we filter out the classes with fewer than 50 nodes. The transformed dataset is dubbed as CoraFull (Zargarbashi et al., 2023). Detailed statistics of datasets are shown in Appendix E.
Baselines. Since our SNAPS is a general post-processing methods for GNNs, here we choose GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018) and APPNP (Gasteiger et al., 2018) as structure-aware models and MLP as a structure-independent model. Moreover, our SNAPS can be based on general conformal prediction non-conformity scores, here we choose APS (Romano et al., 2020) and RAPS (Angelopoulos et al., 2021). For comparison, we compare not only with the basic scores, i.e., APS and RAPS, but also with DAPS (Zargarbashi et al., 2023) for GNNs.
CP Settings. For the basic model GCN, GAT, APPNP and MLP, we follow parameters suggested by (Zargarbashi et al., 2023). For DAPS, we follow the official implementation. Since GNNs are sensitive to splits, especially in the sparsely labeled setting (Shchur et al., 2019), we train the model over ten trials using varying train/validation splits. For per class in the training/validation set, we randomly select 20 nodes. For Arxiv and Products dataset, we follow the official split in PyTorch Geometric (Fey and Lenssen, 2019). Then, the remaining nodes are included in the calibration set and the test set. The calibration set ratio is suggested by (Huang et al., 2023b), i.e., modifying the calibration set size to $|\mathcal{V}_{\text{calib}}|=\min\{1000,|\mathcal{V}_{\text{calib}}\cup\mathcal{V}_{\text{test}}|/2\}$. For each trained model, we conduct 100 random splits of calibration/test set. Thus, we totally conduct 1000 trials to evaluate the effectiveness of CP. For the non-conformity score function that requires hyper-parameters, we split the calibration set into two sets, one for tuning parameters, and the other for conformal calibration (Zargarbashi et al., 2023). For SNAPS, we choose $\lambda$ and $\mu$ in increments of 0.05 within the range 0 to 1, and ensure that $\lambda+\mu<=1$. Each experiment is done with a single NVIDIA V100 32GB GPU.

### v2

@@ Abstract
Graph Neural Networks have achieved remarkable accuracy in semi-supervised node classification tasks. However, these results lack reliable uncertainty estimates. Conformal prediction methods provide a theoretical guarantee for node classification tasks, ensuring that the conformal prediction set contains the ground-truth label with a desired probability (e.g., 95%). In this paper, we empirically show that for each node, aggregating the non-conformity scores of nodes with the same label can improve the efficiency of conformal prediction sets while maintaining valid marginal coverage. This observation motivates us to propose a novel algorithm named Similarity-Navigated Adaptive Prediction Sets (SNAPS), which aggregates the non-conformity scores based on feature similarity and structural neighborhood. The key idea behind SNAPS is that nodes with high feature similarity or direct connections tend to have the same label. By incorporating adaptive similar nodes information, SNAPS can generate compact prediction sets and increase the singleton hit ratio (correct prediction sets of size one). Moreover, we theoretically provide a finite-sample coverage guarantee of SNAPS. Extensive experiments demonstrate the superiority of SNAPS, improving the efficiency of prediction sets and singleton hit ratio while maintaining valid coverage.
@@ 1 Introduction
Graph Neural Networks (GNNs), which process graph-structured data by the message-passing manner (Kipf and Welling, 2017; Hamilton et al., 2017; Velickovic et al., 2018; Xu et al., 2019), have achieved remarkable accuracy in various high-stakes applications, e.g., drug discovery (Li et al., 2022), fraud detection (Liu et al., 2023) and traffic forecasting (Jiang and Luo, 2022), where any erroneous prediction can be costly and dangerous (Amodei et al., 2016; Gao et al., 2019). To improve the reliability of prediction results, many methods have been investigated to quantify the model uncertainty (Gal and Ghahramani, 2016; Guo et al., 2017; Kendall and Gal, 2017; Wang et al., 2021; Hsu et al., 2022; Tang et al., 2024), while these methods lack theoretical guarantees of quantification. Conformal prediction (CP), on the other hand, offers a systematic approach to construct prediction sets that contain ground-truth labels with a desired coverage guarantee  (Vovk et al., 2005; Romano et al., 2020; Angelopoulos et al., 2021; Huang et al., 2023a; Xi et al., 2024).
@@
CP algorithms utilize non-conformity scores to measure dissimilarity between a new instance and the training instances. The lower the score of a new instance, the more likely it belongs to the same distribution space as the training instances, thereby included in the prediction set. To improve the efficiency of prediction sets for GNNs, DAPS (Zargarbashi et al., 2023) smooths node-wise non-conformity scores by incorporating neighborhood information based on the assumption of network homophily. Similar to DAPS, CF-GNN (Huang et al., 2023b) introduces a topology-aware output correction model that learns to update prediction and then produces more efficient prediction sets or intervals with the inefficiency as the optimization objective. However, they only consider structural neighbors and ignore the effect of other nodes that are far from the ego node. This motivates us to analyze the influence of global nodes on the size of prediction sets.
@@
In this work, we show that aggregating the information of global nodes with the same label as the ego node benefits the performance of CP methods. We provide an empirical analysis by randomly selecting nodes with the same label as the ego node from an oracle perspective, where the ground-truth labels of all nodes are known, and then aggregating their non-conformity scores into the ego node. The results indicate that aggregating scores of these nodes can significantly reduce the average size of prediction sets. This suggests that the information of nodes with the same label could correct the non-conformity scores, thereby prompting the efficiency of prediction sets. Detailed analysis is presented in Subsection 3.1. However, during the testing phase, the ground-truth label of every test node is unknown. Inspired by the analysis, our key idea is to accurately identify and select as many nodes with the same label as the ego node as possible and aggregate their non-conformity scores.
@@
To verify the effectiveness of our method, we conduct thorough empirical evaluations on 10 datasets, including both small datasets and large-scale datasets, e.g., OGBN Products (Bhatia et al., 2016). The results demonstrate that SNAPS not only achieves the pre-defined empirical marginal coverage but also achieves better performance over the compared methods. For example, on OGBN Products, our method reduces the average size of prediction sets from 14.92 of APS to 7.68. Moreover, we adapt SNAPS to image classification problems. The results demonstrate that SNAPS reduces the average size of prediction sets from 19.639 to 4.079 – only $\frac{1}{5}$ of the prediction set size from APS on ImageNet (Deng et al., 2009). Code is available at https://github.com/janqsong/SNAPS.
@@
  • We empirically explain that non-conformity scores of nodes with the same label as the ego node play a critical role in their non-conformity scores.
@@
  • Extensive experimental results demonstrate the effectiveness of our proposed method. We show that SNAPS not only maintains the pre-defined coverage but also achieves great performance in efficiency and singleton hit ratio.
@@ 2 Preliminary
Adaptive Prediction Sets. In the APS method, the non-conformity scores are calculated by accumulating the softmax probabilities in descending order. Formally, given a data pair $(\boldsymbol{x},y)$ and a predicted probability estimator $\pi(\boldsymbol{x})_{y}$ for $(\boldsymbol{x},y)$, where $\pi(\boldsymbol{x})_{y}$ is the predicted probability for class $y$, the non-conformity scores can be computed by:
$s(\boldsymbol{x},y)=\sum_{i=1}^{|\mathcal{Y}|}\pi(\boldsymbol{x})_{i}\mathbb{I}[\pi(\boldsymbol{x})_{i}>\pi(\boldsymbol{x})_{y}]+\xi\cdot\pi(\boldsymbol{x})_{y},$ (3)
where $\xi\in[0,1]$ is a uniformly distributed random variable. Then, the prediction set is constructed as $\mathcal{C}(\boldsymbol{x})=\{y|s(\boldsymbol{x},y)\leq\hat{q}\}$.
Evaluation Metrics. The goal is to improve the efficiency of conformal prediction sets as much as possible while maintaining the empirical marginal coverage guarantee. Given the testing nodes set $\mathcal{V}_{\text{test}}$, the efficiency is defined as the average size of prediction sets: $\mathrm{Size}:=\frac{1}{|\mathcal{V}_{\text{test}}|}\sum_{v_{i}\in\mathcal{V}_{\text{test}}}|\mathcal{C}(\boldsymbol{x}_{i})|.$ The smaller the size, the more efficient CP is. The empirical marginal coverage is defined as $\mathrm{Coverage}:=\frac{1}{|\mathcal{V}_{\text{test}}|}\sum_{v_{i}\in\mathcal{V}_{\text{test}}}\mathbb{I}[y_{i}\in\mathcal{C}(\boldsymbol{x}_{i})]$. Although efficiency is a common metric for evaluating CP, singleton hit ratio (SH), defined as the proportion of prediction sets of size one that contains the ground-truth label, is also important (Zargarbashi et al., 2023). The formula of SH is defined as: $\mathrm{SH}:=\frac{1}{|\mathcal{V}_{\text{test}}|}\sum_{v_{i}\in\mathcal{V}_{\text{test}}}\mathbb{I}[\mathcal{C}(\boldsymbol{x}_{i})=\{y_{i}\}]$.
@@ 3 Motivation and Methodology
In this section, we begin by outlining our motivation, substantiating its validity and feasibility through experimental evidence. Then, we propose our method, SNAPS. Finally, we demonstrate that SNAPS satisfies the exchangeability assumption required by CP and offer proof of its improved efficiency compared to basic non-conformity score methods.
@@ 3.2 Similarity-Navigated Adaptive Prediction Sets
$\mathrm{Sim}(i,j)=\frac{\boldsymbol{x}_{i}^{\top}\boldsymbol{x}_{j}}{\|\boldsymbol{x}_{i}\|_{2}\cdot\|\boldsymbol{x}_{j}\|_{2}},$ (4)
@@
Figure 3: The overall framework of SNAPS. (1) Basic non-conformity score function. We first use basic non-conformity score functions, e.g., APS, to convert node embeddings into non-conformity scores. (2) SNAPS function. We then aggregate basic non-conformity scores of $k$-NN with feature similarity and one-hop structural neighbors to correct the non-conformity scores of nodes. (3) Conformal Prediction. Finally, we use conformal prediction to generate prediction sets, significantly reducing their size compared to the basic score functions.
@@
where $s(\cdot,\cdot)$ is the basic non-conformity score function and $\hat{s}(\cdot,\cdot)$ is the SNAPS score function. Both $\lambda$ and $\mu$ are hyperparameters, which are used to measure the importance of three parts of non-conformity scores. The framework of SNAPS is shown in Figure 3 and the pseudo-code is in Appendix B.
@@ 3.3 Theoretical Analysis
To deploy CP for graph-structured data, the only assumption we should satisfy is exchangeability, i.e., the joint distribution of calibration and testing sets remains unchanged under any permutation. Several studies have demonstrated that non-conformity scores based on the node embeddings obtained by any GNN models are invariant to the permutation of nodes while permuting their edges correspondingly in the calibration and testing sets. This invariance arises because GNNs models and non-conformity score functions only use the structures and attributes in the graph, without dependence on the order of the nodes (Zargarbashi et al., 2023; Huang et al., 2023b). Under this condition, we prove that SNAPS non-conformity scores are still exchangeable.
@@
Assume that all of the nodes aggregated by SNAPS are the same label as the ego node. Given a data pair $(\boldsymbol{x},y)$ and a predicted estimator $\pi(\boldsymbol{x})_{y}$ for $(\boldsymbol{x},y)$, where $\pi(\boldsymbol{x})_{y}$ is the predicted probability for class $y$. Moreover, $\epsilon_{ki}$ reflects the model’s error in misclassifying the ground-truth label $k$ as label $i$. Let $\boldsymbol{S}$ be APS scores of nodes, where $\boldsymbol{S}_{ui}\in[0,1]$ is the score corresponding to node $u$ with label $i$. Let $E_{k}[\pi(\boldsymbol{x}_{u})_{i}]$ and $E_{k}[\boldsymbol{S}_{ui}]$ be the average of predicted probability and scores corresponding to label $i$ of nodes whose ground-truth labels are $k$, respectively. Let $\eta$ be $1-\alpha$ quantile of basic non-conformity scores with a significance level $\alpha$. If $E_{k}[\boldsymbol{S}_{uk}]<\eta$ and $E_{k}[\boldsymbol{S}_{ui}]\geq(1-\epsilon_{ki})E_{k}[\pi(\boldsymbol{x}_{u})_{max}]+E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]$, where $E_{k}[\pi(\boldsymbol{x}_{u})_{max}]$ denotes the maximum predicted probability of nodes whose ground-truth labels are $k$ and $\xi\in[0,1]$ is a uniformly distributed random variable, then
$\mathbb{E}[|\mathcal{\tilde{C}}(\boldsymbol{x})|]\leq\mathbb{E}[|\mathcal{C}(\boldsymbol{x})|],$
where $\mathcal{C}(\cdot)$ and $\mathcal{\tilde{C}}(\cdot)$ represent the prediction set from the APS score function and SNAPS score function, respectively.
In other words, SNAPS consistently generates a smaller prediction set than basic non-conformity scores functions and maintains the desired marginal coverage rate. It is obvious that we can’t ignore a very important thing, which is to select nodes with the same label as the ego node as correctly as possible, otherwise, it will lead to a decrease in the efficiency of SNAPS.
@@ 4 Experiments
In this section, we conduct extensive experiments on semi-supervised node classification to demonstrate the effectiveness of SNAPS on graph-structure data. We also adapt SNAPS for image classification problems. Furthermore, we perform ablation studies and parameter analysis to show the importance of different components in SNAPS and evaluate its robustness, respectively.
@@ 4.1 Experimental Settings
Datasets. In our experiments, we consider ten datasets with high homophily, where connected nodes in the graph are likely to share the same label. These datasets include the common citation graphs: CoraML (McCallum et al., 2000), PubMed (Namata et al., 2012), CiteSeer (Sen et al., 2008), CoraFull (Bojchevski and Günnemann, 2018), Coauthor Physics (Physics) and Coauthor CS (CS) (Shchur et al., 2018) and the co-purchase graphs: Amazon Photos (Photos) and Amazon Computers (Computers) (McAuley et al., 2015; Shchur et al., 2018). Moreover, we consider two large-scale graph datasets, i.e., OGBN Arxiv (Arxiv) (Wang et al., 2020) and OGBN Products (Products) (Bhatia et al., 2016). Particularly, for CoraFull which is highly class-imbalanced, we filter out the classes with fewer than 50 nodes. The transformed dataset is dubbed as CoraFull (Zargarbashi et al., 2023). Detailed statistics of these datasets are shown in Appendix F. In addition to the datasets mentioned above, we discuss two heterophilous graph datasets in Appendix C.1, namely Chameleon and Squirrel, both of which are two Wikipedia networks (Rozemberczki et al., 2021).
Baselines. Since our SNAPS is a general post-processing method for GNNs, here we choose GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018) and APPNP (Gasteiger et al., 2018) as structure-aware models and MLP as a structure-independent model. Moreover, our SNAPS can be based on general conformal prediction non-conformity scores, here we choose APS (Romano et al., 2020) and RAPS (Angelopoulos et al., 2021). For comparison, we compare not only with the basic scores, i.e., APS and RAPS, but also with DAPS (Zargarbashi et al., 2023) for GNNs.
CP Settings. For the basic model GCN, GAT, APPNP and MLP, we follow parameters suggested by (Zargarbashi et al., 2023). For DAPS, we follow the official implementation. Since GNNs are sensitive to splits, especially in the sparsely labeled setting (Shchur et al., 2018), we train the model over ten trials using varying train/validation splits. For per class in the training/validation set, we randomly select 20 nodes. For Arxiv and Products dataset, we follow the official split in PyTorch Geometric (Fey and Lenssen, 2019). Then, the remaining nodes are included in the calibration set and the test set. The calibration set ratio is suggested by (Huang et al., 2023b), i.e., modifying the calibration set size to $|\mathcal{V}_{\text{calib}}|=\min\{1000,|\mathcal{V}_{\text{calib}}\cup\mathcal{V}_{\text{test}}|/2\}$. For each trained model, we conduct 100 random splits of calibration/test set. Thus, we totally conduct 1000 trials to evaluate the effectiveness of CP. For the non-conformity score function that requires hyper-parameters, we split the calibration set into two sets, one for tuning parameters, and the other for conformal calibration (Zargarbashi et al., 2023). For SNAPS, we choose $\lambda$ and $\mu$ in increments of 0.05 within the range 0 to 1, and ensure that $\lambda+\mu<=1$. Each experiment is done with a single NVIDIA V100 32GB GPU.